OpenStudy (anonymous):

Express answer in exact form. Find the area of the larger segment whose chord is 8" long in a circle with an 8" radius.

OpenStudy (p0sitr0n):

proportion. C= 2 pi r Chord = 8/16pi A = pir^2 Achord =x 1/2pi * pi r^2 = Achord 1/2 * r^2 = 64/2 = 32

OpenStudy (anonymous):

Thanks!